OpenStudy (anonymous):

HELP HARD MATH QUESTION!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (anonymous):

Nakim simplified \[3\sqrt{2X} + X \sqrt{8X} - 5 \sqrt{18} and got -10X \sqrt{2X} \] for an answer

OpenStudy (anonymous):

Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

jhonyy9 (jhonyy9):

what is your opinion from this ? can you solve it right here ?

OpenStudy (anonymous):

no lol i need help :(

jhonyy9 (jhonyy9):

so because the first term is 3sqrt(2x) you need simplifie every terms in the same station - so than you need to simplifie xsqrt(8x) = ? so and from -5sqrt(18) a term simplified on what you can do it ok ?

jhonyy9 (jhonyy9):

so sqrt(8x) = what is 8 using exponent,power of 2 ?

OpenStudy (anonymous):

4?

jhonyy9 (jhonyy9):

this sign that 2 on exponent 4 equal 8 not is right

OpenStudy (anonymous):

huh?

jhonyy9 (jhonyy9):

so if 8=2*2*2 than result that 2 on exponent ??? equal 8

OpenStudy (anonymous):

3 :D

jhonyy9 (jhonyy9):

yes - so this is very easy yes ?

OpenStudy (anonymous):

wait can u explain step by step how to SOLVE THIS !!!! Whats the first step and use the equation tab lol

jhonyy9 (jhonyy9):

than xsqrt(8x) = ?

jhonyy9 (jhonyy9):

yes i wann to make you solving this step by step here now

OpenStudy (anonymous):

okay so we need simply everything

jhonyy9 (jhonyy9):

yes than just you need answering my questions please ok ?

OpenStudy (anonymous):

\[3\sqrt{2X}\] okay can we simplify that?

jhonyy9 (jhonyy9):

but first there is xsqrt8x =

OpenStudy (anonymous):

we do that one next

jhonyy9 (jhonyy9):

yes because 3sqrt2x not can simplifie again

OpenStudy (anonymous):

oh so that one is simplified okay so now we do \[X \sqrt{8X}\]

jhonyy9 (jhonyy9):

ys how can you simplifie it using this sqrt8 = sqrt(2*2*2) =

OpenStudy (anonymous):

\[\sqrt{2 * 2 * 2 }\]

jhonyy9 (jhonyy9):

yes how you simplifie it ?

OpenStudy (anonymous):

hmmm idk show me please?

jhonyy9 (jhonyy9):

ok sqrt(2*2*2) = sqrt(2^2 *2) = sqrt2^2 *sqrt2 so than sqrt2^2 =

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

hmmm not surel ol

jhonyy9 (jhonyy9):

why how many is sqrt(x^2) = ???

OpenStudy (anonymous):

2*2

jhonyy9 (jhonyy9):

what is 2*2 ?

OpenStudy (anonymous):

idk im confused? LOL

jhonyy9 (jhonyy9):

so if do you know that sqrt(x^2) =x so than sqrt(2^2) = ?

OpenStudy (anonymous):

X :D

OpenStudy (anonymous):

right?

jhonyy9 (jhonyy9):

not is so if sqrt(x^2) = x so than sqrt(2^2) =2 ok ?

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

so than sqrt(2^3) = ?

OpenStudy (anonymous):

3?

jhonyy9 (jhonyy9):

when you know that sqrt(2^2)=2 and sqrt(2^3) =sqrt(2^2 *2) so than how you have got 3 ?

jhonyy9 (jhonyy9):

sqrt(2^3) = sqrt(2^2 *2)=sqrt(2^2) *sqrt2 =

jhonyy9 (jhonyy9):

there is all what you need to know check it please

OpenStudy (anonymous):

so what does it look like simplified?

jhonyy9 (jhonyy9):

so if sqrt2^2 =2 than result that sqrt(2^2 *2) = 2sqrt2 yes ?

jhonyy9 (jhonyy9):

so than xsqrt8x = ?

OpenStudy (anonymous):

So it looks like this right \[X \sqrt{8X} = 2\sqrt{2}\]

jhonyy9 (jhonyy9):

so and where is x ?

OpenStudy (anonymous):

hmmm on top of the 2 on the inside right?

jhonyy9 (jhonyy9):

write it please !

OpenStudy (anonymous):

\[X \sqrt{8X} = 2\sqrt{2^{^{2}}}\]

OpenStudy (anonymous):

i meant to put X lol

jhonyy9 (jhonyy9):

how ?

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

what would it look like?

jhonyy9 (jhonyy9):

xsqrt8x = ?

OpenStudy (anonymous):

idk

jhonyy9 (jhonyy9):

xsqrt8x = xsqrtx * sqrt8 = ?

OpenStudy (anonymous):

?

jhonyy9 (jhonyy9):

do you know that sqrt8 = ?

OpenStudy (anonymous):

no

jhonyy9 (jhonyy9):

before you have solved it

OpenStudy (anonymous):

im confuised

jhonyy9 (jhonyy9):

look it little up

OpenStudy (anonymous):

k

OpenStudy (anonymous):

just tell me

jhonyy9 (jhonyy9):

you need to learn it how you can solve it ,right ?

OpenStudy (anonymous):

just tell me this part so we can move on

jhonyy9 (jhonyy9):

ok look so if sqrt8 = 2sqrt2 so than xsqrt8x = 2xsqrt(2x) do you know how i have got it ?

OpenStudy (anonymous):

okay so thats the simplyfy right?

jhonyy9 (jhonyy9):

yes but you need to continue how is your exercise again ?

OpenStudy (anonymous):

so the answer is \[X \sqrt{8X} = 2X \sqrt{2X}\]

jhonyy9 (jhonyy9):

for this midle term yes

OpenStudy (anonymous):

okay lol and the last one to simply is \[-5\sqrt{18X}\]

jhonyy9 (jhonyy9):

but in your exercise you not have wrote it hence ,check it please

jhonyy9 (jhonyy9):

there is -5sqrt18 right ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so than how is correct ?

OpenStudy (anonymous):

what do u mean? thats the next part lol we have to do

jhonyy9 (jhonyy9):

so check it please your exercise how you have wrote firstly and how you have wrote it here little up

jhonyy9 (jhonyy9):

inside radical need being x or not ?

OpenStudy (anonymous):

oh i forgot to put x at the beggind its suppose to be x

jhonyy9 (jhonyy9):

so than correct is -5sqrt18x yes ,

jhonyy9 (jhonyy9):

so than how you can simplifie it ?

OpenStudy (anonymous):

hmmm im not sure

jhonyy9 (jhonyy9):

why if do you know that 18=2*9=2*3^2 so than how will be simplified ?

OpenStudy (anonymous):

2sqrt9

jhonyy9 (jhonyy9):

why ? when sqrt9 = sqrt3^2 =3

OpenStudy (anonymous):

hmm true so 7sqrt2

jhonyy9 (jhonyy9):

how you have got 7 ?

OpenStudy (anonymous):

im not sure lol

jhonyy9 (jhonyy9):

so there is -5sqrt18x =

jhonyy9 (jhonyy9):

when sqrt18=sqrt(3^2 *2) =3sqrt2

OpenStudy (anonymous):

hmm okay

jhonyy9 (jhonyy9):

so than -5sqrt18x = ?

jhonyy9 (jhonyy9):

-15sqrt2x why ?

OpenStudy (anonymous):

what does it =?

jhonyy9 (jhonyy9):

i wann to saying you what is the right answer for